A is 8% taller than B, while B is 9% taller than C. by how much

in percent is A taller than C?


1
Expert's answer
2021-09-22T00:52:42-0400

Let

"A =[\frac{100+8}{100}]B=1.08B"


"B =[\frac{100+9}{100}]C=1.09C"



"A=(1.08)1.09C=1.1772C"


Hence "(1.1772\u00d7100)-100=17.72"


A taller than C by 17.72%